# Artifacts Are Not Noise: Embodied Resonance and the 70% Signal Loss in Conventional EEG

### v2

@@ Abstract
Current AI systems excel at pattern recognition but fail at causal reasoning. We argue this is not an engineering limitation but reveals something fundamental about the nature of understanding itself. We propose that causal cognition requires a specific physical architecture: stochastic, coupled oscillators with whole-system coordination. To test this, we analyzed high-density EEG (64 channels, 10 subjects, 500+ trials) from a P300 target recognition task. We computed the Kuramoto Order Parameter ($R$) to measure global phase synchronization and compared it to standard voltage (ERP) and coherence (ITC) metrics. Four findings establish the framework. Phase and voltage are globally independent ($r=0.048$) yet strongly trial-coupled ($r=0.590$), proving $R$ captures hidden cognitive structure. Voltage precedes phase by 293 ms, revealing sequential computation. Frequency decomposition shows Theta (169 ms), Alpha (286 ms), and Beta (777 ms) cascade. Our metric is distinct from standard ITC ($r=0.155$). Then we tested the standard assumption. Conventional artifact rejection removes eye movements, muscle activity, and autonomic signals before analysis. This assumes:
$\text{Cognition}=\text{Neural Activity}+\text{Noise}$
We ran the identical analysis with and without rejection. Removing artifacts reduced the trial-level correlation threefold ($0.590\to 0.195$). Target discrimination reversed sign ($0.6\%\to-0.4\%$). What we discard as noise is 70% of the signal. This falsifies the equation above. Cognition is not isolated neural computation. It is whole-body phase synchronization spanning neural, muscular, and autonomic systems. For AI, the implications are direct: embodied sensorimotor integration is not optional. It is the substrate that makes understanding possible.
1 Introduction
Artificial intelligence systems can interpolate patterns but cannot reason about causality [1, 2, 3]. This is the "Kepler versus Newton" problem: current AI describes what happens but cannot explain why. We argue this limitation is architectural, not addressable through scale or training. The issue is substrate. Understanding requires four components working in closed loop: stochastic exploration, bounded agency, intrinsic cost, and iterative feedback. Deterministic digital systems lack the first. Disembodied systems lack the second. Most critically, they lack the physical dynamics that convert noise into structured proposals. We solve this with the Resonance Principle. Model the cognitive substrate as coupled oscillators [4]. Noise is not the idea. It perturbs the system. Ideas emerge as stable resonant modes. Learning sculpts this landscape through cost feedback, stabilizing useful patterns and destabilizing others. If this is correct, human cognition should show physical signatures of resonance during tasks requiring causal discrimination. We test this using P300 target recognition, where subjects distinguish targets from non-targets. This is foundational causal inference: determining whether a stimulus is self-relevant or random.
@@ 1.1 Testing the Artifact Assumption
But if cognition requires whole-organism coordination, this equation is wrong. Eye movements might be part of attention. Muscle tension might encode arousal. Autonomic signals might participate in decision-making. We test Equation 1 directly. If artifacts are noise, removing them should strengthen cognitive signals. If they contain information, removal should weaken signals. Prediction: Standard preprocessing destroys rather than preserves cognitive structure.
2 Methods
@@ 2.2.2 Artifact Rejection Comparison
Conventional (Clean): Epochs exceeding 150 µV peak-to-peak amplitude in any channel were rejected (standard threshold for artifacts). Average rejection rate: 0.60%.
Whole-System (Raw): Identical pipeline without rejection, preserving all trials including those with eye movements, muscle activity, and autonomic signals. This tests Equation 1. If artifacts are noise, Clean should outperform Raw. If artifacts contain signal, Raw should outperform Clean.
@@ 2.5 Statistical Analysis
    Global: Correlation between grand-average $R(t)$ and $ERP(t)$ time series
@@
    Trial-level: Correlation between peak $R$ and peak absolute $ERP$ within each trial (100-600 ms window)
@@
    Temporal: Cross-correlation to identify lag between $R(t)$ and $ERP(t)$
@@
    Novelty: Correlation between $R$ and standard ITC
@@ 3.6 Artifact Rejection Destroys the Signal
This result falsifies Equation 1. If artifacts were noise, their removal would strengthen rather than weaken cognitive signals. The data demonstrate that signals conventionally discarded contain approximately 70% of the task-relevant variance.
@@ 3.7 Causal Mediation Analysis
Dose-Response. While limited by data sparsity in high-amplitude bins (only low-amplitude trials: $r=0.002\pm 0.054$), the intervention effect itself (Table 1) constitutes the strongest dose-response evidence: complete artifact removal yields threefold signal reduction, demonstrating causal relationship between artifact presence and cognitive coupling strength. These five tests, combined with the intervention result, provide six converging lines of causal evidence. The pattern—regional specificity, temporal integration, confound resistance, dynamic coupling, and dose-dependent effects—is inconsistent with artifacts as random noise contamination. Instead, it supports artifacts as integral components of whole-body cognitive computation.
@@ 4.2 The Central Finding
Conventional artifact rejection reduced the cognitive signal threefold. This falsifies the foundational assumption of modern EEG preprocessing. Eye movements, muscle activity, and autonomic signals are not noise. They contain most of the predictive variance. Causal mediation is supported by six independent lines of evidence. Beyond the intervention effect (Table 1), we demonstrate regional specificity (frontal > temporal > occipital), temporal integration (near-synchronous coupling), confound control (survives baseline correction), within-trial coupling (moment-to-moment covariation), and dose-dependence (complete removal yields threefold loss). This convergence approaches the strongest causal inference possible from observational data [1]. Three observations support this interpretation. First, if artifacts were random noise, they would corrupt all conditions equally. Instead, their removal specifically eliminated target versus non-target discrimination, proving task-relevant information. Second, the threefold correlation drop indicates 70% of variance resides in what we discard. Third, temporal structure (293 ms lag, frequency cascade) remained consistent regardless of preprocessing, validating genuine physiological process rather than artifact contamination. We propose:
@@ 4.4 Implications for EEG Research
If correct, decades of research may have systematically discarded the signals they sought. The P300, studied as pure cognitive component after artifact rejection, might be downstream effect of whole-body cascade that requires eye movements, postural shifts, and autonomic activation. This does not invalidate prior work. It recontextualizes it. Isolated neural signals contain information, just 70% less than whole-system signals. Future research should reconsider whether artifacts represent noise or data.
@@ 4.5 Implications for Artificial Intelligence
This provides empirical support for embodied AI as computational necessity rather than philosophical preference. The physicality of biological systems (intrinsic noise, analog coupling, sensorimotor feedback) is not incidental. It is essential to the resonance dynamics that enable understanding. Digital AI lacking body resonance is constrained like artifact-rejected EEG. This validates neuromorphic approaches: systems with stochastic dynamics, analog coupling, and embodied sensorimotor loops. Not to simulate brains, but to instantiate the physics.
@@ 4.6 Limitations
This study has limitations. Sample size is modest (10 subjects), though 500+ trials provide statistical power for within-subject effects. Single task type (P300) requires replication across domains. Correlational design cannot establish definitive causation despite six converging tests; perturbation studies would strengthen claims. Absence of behavioral measures (reaction time, accuracy) limits functional interpretation. The dose-response analysis was limited by data sparsity in high-amplitude bins. Nevertheless, the threefold signal loss from artifact rejection is quantitative, replicable, and challenges foundational assumptions. The finding invites reconsideration of what constitutes "signal" versus "noise" in cognitive neuroscience.
5 Conclusion
The Kuramoto Order Parameter reveals cognitive structure invisible to standard voltage analysis. Phase synchronization is globally independent of voltage ($r=0.048$) yet trial-level predictive ($r=0.590$). Initial voltage precedes phase by 293 ms. Frequency decomposition shows Theta, Alpha, Beta cascade. The metric is distinct from standard coherence measures ($r=0.155$). Conventional artifact rejection eliminates 70% of the cognitive signal. What neuroscience discards as noise—eye movements, muscle activity, autonomic responses—contains most predictive variance. Six converging lines of causal evidence (intervention, regional specificity, temporal integration, confound control, within-trial coupling, dose-dependence) support artifacts as integral computational components rather than random contamination. This falsifies the assumption that cognition equals isolated neural computation. Understanding emerges from whole-body resonance. For artificial intelligence, disembodied digital architectures may be fundamentally constrained. Neuromorphic systems implementing stochastic, embodied dynamics may be necessary to achieve genuine causal reasoning.
Acknowledgments
The author thanks Professor Ehab Emam for invaluable guidance and feedback on this manuscript. Data were obtained from PhysioNet (doi:10.13026/C2101S). All analysis code will be made available upon publication.

### v3

@@ Abstract
Current AI systems excel at pattern recognition but fail at causal reasoning. We argue this is not an engineering limitation but reveals something fundamental about the nature of understanding itself. We propose that causal cognition requires a specific physical architecture: stochastic, coupled oscillators with whole-system coordination. To test this, we analyzed high-density EEG (64 channels, 10 subjects, 500+ trials) from a P300 target recognition task. We computed the Kuramoto Order Parameter ($R$) to measure global phase synchronization and compared it to standard voltage (ERP) and coherence (ITC) metrics. Four findings establish the framework. Phase and voltage are globally independent ($r=0.048$) yet strongly trial-coupled ($r=0.590$), proving $R$ captures hidden cognitive structure. Voltage precedes phase by 293 ms, revealing sequential computation. Frequency decomposition shows Theta (169 ms), Alpha (286 ms), and Beta (777 ms) cascade. Our metric is distinct from standard ITC ($r=0.155$). Then we tested the standard assumption. Conventional artifact rejection removes eye movements, muscle activity, and autonomic signals before analysis. This assumes Cognition = Neural Activity + Noise. We ran the identical analysis with and without rejection. Removing artifacts reduced the trial-level correlation threefold ($0.590\to 0.195$). Target discrimination reversed sign ($0.6\%\to-0.4\%$). What we discard as noise is 70% of the signal. This falsifies the equation above. Cognition is not isolated neural computation. It is whole-body phase synchronization spanning neural, muscular, and autonomic systems. For AI, the implications are direct: embodied sensorimotor integration is not optional. It is the substrate that makes understanding possible.
Keywords: Embodied Cognition, EEG Artifacts, Kuramoto Order Parameter, P300, Neuromorphic Computing, Causal Reasoning
  ntroduction
Artificial intelligence systems can interpolate patterns but cannot reason about causality [1, 2, 3]. This is the "Kepler versus Newton" problem: current AI describes what happens but cannot explain why. We argue this limitation is architectural, not addressable through scale or training.
The issue is substrate. Understanding requires four components working in closed loop: stochastic exploration, bounded agency, intrinsic cost, and iterative feedback. Deterministic digital systems lack the first. Disembodied systems lack the second. Most critically, they lack the physical dynamics that convert noise into structured proposals.
We solve this with the Resonance Principle. Model the cognitive substrate as coupled oscillators [4]. Noise is not the idea. It perturbs the system. Ideas emerge as stable resonant modes. Learning sculpts this landscape through cost feedback, stabilizing useful patterns and destabilizing others.
If this is correct, human cognition should show physical signatures of resonance during tasks requiring causal discrimination. We test this using P300 target recognition, where subjects distinguish targets from non-targets. This is foundational causal inference: determining whether a stimulus is self-relevant or random.
@@ 1.1  Testing the Artifact Assumption
But if cognition requires whole-organism coordination, this equation is wrong. Eye movements might be part of attention. Muscle tension might encode arousal. Autonomic signals might participate in decision-making.
We test Equation 1 directly. If artifacts are noise, removing them should strengthen cognitive signals. If they contain information, removal should weaken signals.
Prediction: Standard preprocessing destroys rather than preserves cognitive structure.
  ethods
@@ 2.2.2 Artifact Rejection Comparison
Conventional (Clean): Artifacts were removed using Independent Component Analysis (ICA). EOG and muscle components were identified via correlation with frontal electrodes and surgically subtracted from the data. This preserves the neural signal while removing the artifact, serving as a stricter test of the ’noise’ hypothesis.
Whole-System (Raw): Identical pipeline without rejection, preserving all trials including those with eye movements, muscle activity, and autonomic signals.
This tests Equation 1. If artifacts are noise, Clean should outperform Raw. If artifacts contain signal, Raw should outperform Clean.
@@ 2.5  Statistical Analysis
    Global: Correlation between grand-average $R(t)$ and $ERP(t)$ time series.
@@
    Trial-level: Correlation between peak $R$ and peak absolute $ERP$ within each trial (100-600 ms window).
@@
    Temporal: Cross-correlation to identify lag between $R(t)$ and $ERP(t)$.
@@
    Novelty: Correlation between $R$ and standard ITC.
@@ 3.6  Artifact Rejection Destroys the Signal
This falsifies Equation 1. If artifacts were noise, their removal would strengthen rather than weaken cognitive signals. The data demonstrate that signals conventionally discarded contain approximately 70% of the task-relevant variance.
@@ 3.7  Causal Mediation Analysis
Dose-Response. While limited by data sparsity in high-amplitude bins (only low-amplitude trials: $r=0.002\pm 0.054$), the intervention effect itself (Table 1) constitutes the strongest dose-response evidence: complete artifact removal yields threefold signal reduction, demonstrating causal relationship between artifact presence and cognitive coupling strength.
These five tests, combined with the intervention result, provide six converging lines of causal evidence. The pattern—regional specificity, temporal integration, confound resistance, dynamic coupling, and dose-dependent effects—is inconsistent with artifacts as random noise contamination. Instead, it supports artifacts as integral components of whole-body cognitive computation.
@@ 4.2  The Central Finding
Conventional artifact rejection reduced the cognitive signal threefold. This falsifies the foundational assumption of modern EEG preprocessing. Eye movements, muscle activity, and autonomic signals are not noise. They contain most of the predictive variance.
Causal mediation is supported by six independent lines of evidence. Beyond the intervention effect (Table 1), we demonstrate regional specificity (frontal > temporal > occipital), temporal integration (near-synchronous coupling), confound control (survives baseline correction), within-trial coupling (moment-to-moment covariation), and dose-dependence (complete removal yields threefold loss). This convergence approaches the strongest causal inference possible from observational data [1].
Three observations support this interpretation. First, if artifacts were random noise, they would corrupt all conditions equally. Instead, their removal specifically eliminated target versus non-target discrimination, proving task-relevant information. Second, the threefold correlation drop indicates 70% of variance resides in what we discard. Third, temporal structure (293 ms lag, frequency cascade) remained consistent regardless of preprocessing, validating genuine physiological process rather than artifact contamination.
Critically, if frontal artifacts were merely epiphenomenal markers of attention, their removal should not have eliminated target discrimination (Table 1: $0.6\%\to-0.4\%$). The sign reversal indicates functional contribution, not mere correlation.
We propose:
@@ 4.4  Implications for EEG Research
If correct, decades of research may have systematically discarded the signals they sought. The P300, studied as pure cognitive component after artifact rejection, might be downstream effect of whole-body cascade that requires eye movements, postural shifts, and autonomic activation.
This does not invalidate prior work. It recontextualizes it. Isolated neural signals contain information, just 70% less than whole-system signals. Future research should reconsider whether artifacts represent noise or data.
@@ 4.5  Implications for Artificial Intelligence
This provides empirical support for embodied AI as computational necessity rather than philosophical preference. The physicality of biological systems (intrinsic noise, analog coupling, sensorimotor feedback) is not incidental. It is essential to the resonance dynamics that enable understanding.
Digital AI lacking body resonance is constrained like artifact-rejected EEG. This validates neuromorphic approaches: systems with stochastic dynamics, analog coupling, and embodied sensorimotor loops. Not to simulate brains, but to instantiate the physics.
@@ 4.6  Limitations
This study has limitations. Sample size is modest (10 subjects), though 500+ trials provide statistical power for within-subject effects. Single task type (P300) requires replication across domains. Correlational design cannot establish definitive causation despite six converging tests; perturbation studies would strengthen claims. Absence of behavioral measures (reaction time, accuracy) limits functional interpretation. The dose-response analysis was limited by data sparsity in high-amplitude bins.
Nevertheless, the threefold signal loss from artifact rejection is quantitative, replicable, and challenges foundational assumptions. The finding invites reconsideration of what constitutes "signal" versus "noise" in cognitive neuroscience.
  onclusion
The Kuramoto Order Parameter reveals cognitive structure invisible to standard voltage analysis. Phase synchronization is globally independent of voltage ($r=0.048$) yet trial-level predictive ($r=0.590$). Initial voltage precedes phase by 293 ms. Frequency decomposition shows Theta, Alpha, Beta cascade. The metric is distinct from standard coherence measures ($r=0.155$).
Conventional artifact rejection eliminates 70% of the cognitive signal. What neuroscience discards as noise—eye movements, muscle activity, autonomic responses—contains most predictive variance. Six converging lines of causal evidence (intervention, regional specificity, temporal integration, confound control, within-trial coupling, dose-dependence) support artifacts as integral computational components rather than random contamination. This falsifies the assumption that cognition equals isolated neural computation.
Understanding emerges from whole-body resonance. For artificial intelligence, disembodied digital architectures may be fundamentally constrained. Neuromorphic systems implementing stochastic, embodied dynamics may be necessary to achieve genuine causal reasoning.
ata Availability Statement
Publicly available datasets were analyzed in this study. This data can be found here: PhysioNet (doi:10.13026/C2101S).
thics Statement
Ethical review and approval was not required for the study on human participants in accordance with the local legislation and institutional requirements. Written informed consent for participation was not required for this study in accordance with the national legislation and the institutional requirements.
uthor Contributions
AGE: Conceptualization, Methodology, Software, Formal analysis, Writing—original draft, Visualization.
unding
The author(s) declare that no financial support was received for the research, authorship, and/or publication of this article.
cknowledgments
The author thanks Professor Ehab Emam for invaluable guidance and feedback on this manuscript. Data were obtained from PhysioNet (doi:10.13026/C2101S). The author also acknowledges the preliminary version of this manuscript deposited on arXiv (arXiv:2511.10596). All analysis code will be made available upon publication.
onflict of Interest
The author declares that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.